# A Perfectly Truthful Calibration Measure

### v1

@@ 4 Calibration Errors with Two Bins
In this section, we formally define our new calibration measure: the averaged two-bin calibration error (ATB) as well as its $\ell_{1}$ variant ($\ell_{1}\textup{-}\mathrm{ATB}$). We show that these two errors are quadratically related (Lemma 4.2). We also discuss their various properties: truthfulness (for ATB only), continuity, sample complexity for estimation, and computational efficiency. In the next section (Section 5), we establish their quantitative relationships to smCal and $\underline{\textsc{distCal}}$.
@@
As a simple consequence of Jensen’s inequality, $\ell_{1}\textup{-}\mathrm{ATB}$ is quadratically related to $\mathrm{ATB}$.
@@
Fix a threshold $q$, we write $\Delta_{1}(q)=\operatorname{\mathbb{E}}_{J}[(v-y)\mathbb{I}\left[v<q\right]]$ and $\Delta_{2}(q)=\operatorname{\mathbb{E}}_{J}[(v-y)\mathbb{I}\left[v\geq q\right]]$. The right inequality follows from the fact that $\Delta_{1},\Delta_{2}\in[0,1]$.
@@
$\frac{1}{2}\ell_{1}\textup{-}\mathrm{ATB}(J)^{2}=2\left({\mathbf{E}}_{q}\left[\frac{1}{2}\Delta_{1}(q)+\frac{1}{2}\Delta_{2}(q)\right]\right)^{2}\leq 2{\mathbf{E}}_{q}\left[\frac{1}{2}\Delta_{1}(q)^{2}+\frac{1}{2}\Delta_{2}(q)^{2}\right]=\mathrm{ATB}(J).$
4.1 Truthfulness
From its definition (25), $\mathrm{ATB}$ is clearly a special case of UBSE (Definition 3.1), so its truthfulness follows immediately from Theorem 3.2.
@@ 4.1 Truthfulness
The calibration measure $\mathrm{ATB}$ is truthful on sequences (Definition 2.16). Moreover, $\mathrm{ATB}$ inherits the error decomposition:
@@ 4.3 Sample Complexity
By Proposition B.2, it suffices to prove the following Rademacher complexity bounds for the function families $F=\{f_{q}\}_{q\in[0,1]}$ and $G=\{g_{q}\}_{q\in[0,1]}$ where $f_{q}(v,y)=(v-y)\mathbb{I}\left[v<q\right]$ and $g_{q}(v,y)=(v-y)\mathbb{I}\left[v\geq q\right]$: for every $(v_{1},y_{1}),\ldots,(v_{T},y_{T})\in[0,1]\times\{0,1\}$,
@@
Now consider the family $H=\{h_{q}\}_{q\in[0,1]}$ where $h_{q}(v,y)=\mathbb{I}\left[v<q\right]$. Clearly, $H$ has VC dimension at most $1$. By Proposition B.5, we have
@@
Observe that $f_{q}(v_{i},y_{i})=\eta_{i}(h_{q}(v_{i},y_{i}))$ for every $i=1,\ldots,T$ and $q\in[0,1]$, where $\eta_{i}$ is the uni-variate function $\eta_{i}(u)=(v_{i}-y_{i})u$ for $u\in\mathbb{R}$. Since $|v_{i}-y_{i}|\leq 1$, the function $\eta_{i}$ is $1$-Lipschitz. Therefore, by Proposition B.3, inequality (30) follows from (32). Inequality (31) can be proved similarly. ∎
@@ 4.4 Computational Efficiency
Given $\bm{r},\bm{p}\in[0,1]^{T}$, we can compute $\mathrm{ATB}(\bm{r},\bm{p})$ in $O(T\log T)$ time. Moreover, the algorithm is extremely easy to describe and implement. Define
@@
  • $O(T\log T)$ time: sort predictions in increasing order such that $r_{1}\leq r_{2}\leq\dots\leq r_{T}$.
  • $O(T)$ time: calculate $\Delta_{1}(q)$ for each $q$ in the space of reported predictions by scanning predictions in increasing order.
    Similarly, calculate $\Delta_{2}(q)$ by scanning predictions in decreasing order.
  • $O(T)$ time: Write $r_{0}=0$ and $r_{T+1}=1$. Calculate the expectation over threshold $q$: for each $i=0\dots T$, sum $\Delta_{1}(r_{i})^{2}+\Delta_{2}(r_{i})^{2}$ with weight $|r_{i+1}-r_{i}|$.
A similar algorithm computes $\ell_{1}\textup{-}\mathrm{ATB}(\bm{r},\bm{p})$ in $O(T\log T)$ time. We note that currently known algorithms for computing smCal and $\underline{\textsc{distCal}}$ are much more complicated, with the best known running time being $O(T\log^{2}T)$ and $O(T^{2}\log T)$, respectively (Hu et al., 2024).
5 Approximating the Distance to Calibration Using Two Bins
@@
Corollary 5.2.
Lemma 5.3.
For any distribution $J$ of $(v,y)\in[0,1]\times\{0,1\}$,
Proof.
$\ell_{1}\textup{-}\mathrm{ATB}(\widehat{J})=0.$
By Theorem 4.4,
The lemma is proved by taking the infimum over $\Pi$. ∎
Lemma 5.4.
Proof of Lemma 5.4.
$\frac{1}{T}\sum_{t=1}^{T}(r_{t}-y_{t})w(r_{t})\leq\frac{3}{2}\cdot\mathrm{ATB}(\bm{r},\bm{y}).$ (33)
$=\frac{1}{2}\sum_{s=0}^{T}\Delta_{s}\mathsf{sign}(t-s),$
Summing up over $s=1,\ldots T-1$, we have
$\sum_{s=1}^{T-1}\left|\frac{1}{T}\sum_{t=1}^{T}(r_{t}-y_{t})\Delta_{s}\mathsf{sign}(t-s)\right|\leq\mathrm{ATB}(\bm{r},\bm{y}).$ (35)
Moreover, since $w_{1},w_{T}\in[-1,1]$, we have $|\Delta_{0}|,|\Delta_{T}|\leq 1$. Therefore,
$\left|\frac{1}{T}\sum_{t=1}^{T}(r_{t}-y_{t})\Delta_{0}\mathsf{sign}(t-0)\right|$ $=|\Delta_{0}|\cdot\left|\frac{1}{T}\sum_{t=1}^{T}(r_{t}-y_{t})\right|\leq\mathrm{ATB}(\bm{r},\bm{y}),$ (36)
$\left|\frac{1}{T}\sum_{t=1}^{T}(r_{t}-y_{t})\Delta_{T}\mathsf{sign}(t-T)\right|$ $=|\Delta_{T}|\cdot\left|\frac{1}{T}\sum_{t=1}^{T}(r_{t}-y_{t})\right|\leq\mathrm{ATB}(\bm{r},\bm{y}).$ (37)
Adding up the three inequalities (35) (36) (37) above, we get
$\sum_{s=0}^{T}\left|\frac{1}{T}\sum_{t=1}^{T}(r_{t}-y_{t})\Delta_{s}\mathsf{sign}(t-s)\right|\leq 3\mathrm{ATB}(\bm{r},\bm{y}).$
Proof of Theorem 5.1.
6 Extremely Efficient Optimal Calibration Tester
@@
In this section, we show that our $\mathrm{ATB}$ and $\ell_{1}\textup{-}\mathrm{ATB}$ are both optimally valid for smCal and $\underline{\textsc{distCal}}$, which implies a $O(T\log T)$ time calibration test for smCal and $\underline{\textsc{distCal}}$. It is fairly straightforward to show that $\ell_{1}\textup{-}\mathrm{ATB}$ is $O(1/\sqrt{T})$ valid using its constant approximation to smCal (Theorem 5.1) and its sample complexity bound (Theorem 4.5). In Theorem 6.1 below, we show that $\mathrm{ATB}$ is $O(1/\sqrt{T})$ valid as well, and that this is optimal up to constant.
@@
The proof is completed by choosing $C$ large enough so that $(1/2)(C/3-C^{\prime})^{2}>1$. ∎

### v2

@@ 4 Calibration Errors with Two Bins
In this section, we formally define our calibration measure: the averaged two-bin calibration error (ATB). We show that $\mathrm{ATB}$ satisfies the following properties in the literature: completeness and soundness, truthfulness, continuity, sample complexity for estimation, and computational efficiency. Our proof of the completeness and soundness relies heavily on the quadratic relationship between $\mathrm{ATB}$ and its $\ell_{1}$ variant ($\ell_{1}\textup{-}\mathrm{ATB}$). We will show that $\ell_{1}\textup{-}\mathrm{ATB}$ linearly approximates existing calibration measures, implying the completeness and soundness of both $\ell_{1}\textup{-}\mathrm{ATB}$ and $\mathrm{ATB}$.
@@
To prepare for the proof, $\ell_{1}\textup{-}\mathrm{ATB}$ is quadratically related to $\mathrm{ATB}$ by Jensen’s inequality.
@@
Fix a threshold $q$, we write $\Delta_{1}(q)=\operatorname{\mathbb{E}}_{J}[(v-y)\mathbb{I}\left[v<q\right]]$ and $\Delta_{2}(q)=\operatorname{\mathbb{E}}_{J}[(v-y)\mathbb{I}\left[v\geq q\right]]$. The right inequality follows from the fact that $\Delta_{1},\Delta_{2}\in[-1,1]$.
@@
$\frac{1}{2}\ell_{1}\textup{-}\mathrm{ATB}(J)^{2}=2\left({\operatorname{\mathbb{E}}}_{q}\left[\frac{1}{2}|\Delta_{1}(q)|+\frac{1}{2}|\Delta_{2}(q)|\right]\right)^{2}\leq 2{\operatorname{\mathbb{E}}}_{q}\left[\frac{1}{2}\Delta_{1}(q)^{2}+\frac{1}{2}\Delta_{2}(q)^{2}\right]=\mathrm{ATB}(J).$
4.1 Completeness and Soundness
The completeness and soundness of $\mathrm{ATB}$ follows from the quadratic approximation of existing calibration measure Distance to Calibration.
Theorem 4.3.
Both $\mathrm{ATB}$ and $\ell_{1}\textup{-}\mathrm{ATB}$ are complete and sound.
Section 4.1.1 will prove that $\ell_{1}\textup{-}\mathrm{ATB}$ is a constant approximation of Distance to Calibration. Combined with Lemma 4.2, $\mathrm{ATB}$ is quadratically related to Distance to Calibration, which implies Theorem 4.3.
4.1.1 Approximating the Distance to Calibration Using Two Bins
Theorem 4.4.
Corollary 4.5.
Lemma 4.6.
For any distribution $J$ of $(v,y)\in[0,1]\times\{0,1\}$,
Proof.
$\ell_{1}\textup{-}\mathrm{ATB}(\widehat{J})=0.$
By Theorem 4.9,
The lemma is proved by taking the infimum over $\Pi$. ∎
Lemma 4.7.
Proof of Lemma 4.7.
$\frac{1}{T}\sum_{t=1}^{T}(r_{t}-y_{t})w(r_{t})\leq\frac{3}{2}\cdot\ell_{1}\textup{-}\mathrm{ATB}(\bm{r},\bm{y}).$ (29)
$=\frac{1}{2}\sum_{s=0}^{T}\Delta_{s}\mathsf{sign}(t-s),$
Summing up over $s=1,\ldots T-1$, we have
$\sum_{s=1}^{T-1}\left|\frac{1}{T}\sum_{t=1}^{T}(r_{t}-y_{t})\Delta_{s}\mathsf{sign}(t-s)\right|\leq\ell_{1}\textup{-}\mathrm{ATB}(\bm{r},\bm{y}).$ (31)
Moreover, since $w(r_{1}),w(r_{T})\in[-1,1]$, we have $|\Delta_{0}|,|\Delta_{T}|\leq 1$. Therefore,
$\left|\frac{1}{T}\sum_{t=1}^{T}(r_{t}-y_{t})\Delta_{0}\mathsf{sign}(t-0)\right|$ $=|\Delta_{0}|\cdot\left|\frac{1}{T}\sum_{t=1}^{T}(r_{t}-y_{t})\right|\leq\ell_{1}\textup{-}\mathrm{ATB}(\bm{r},\bm{y}),$ (32)
$\left|\frac{1}{T}\sum_{t=1}^{T}(r_{t}-y_{t})\Delta_{T}\mathsf{sign}(t-T)\right|$ $=|\Delta_{T}|\cdot\left|\frac{1}{T}\sum_{t=1}^{T}(r_{t}-y_{t})\right|\leq\ell_{1}\textup{-}\mathrm{ATB}(\bm{r},\bm{y}).$ (33)
Adding up the three inequalities (31) (32) (33) above, we get
$\sum_{s=0}^{T}\left|\frac{1}{T}\sum_{t=1}^{T}(r_{t}-y_{t})\Delta_{s}\mathsf{sign}(t-s)\right|\leq 3\,\ell_{1}\textup{-}\mathrm{ATB}(\bm{r},\bm{y}).$
Proof of Theorem 4.4.
4.2 (Strict) Truthfulness
From its definition (28), $\mathrm{ATB}$ is clearly a special case of UBSE (Definition 3.1), so its truthfulness follows immediately from Theorem 3.2 and Theorem 3.5.
@@ 4.1.1 Approximating the Distance to Calibration Using Two Bins
The calibration measure $\mathrm{ATB}$ is interim truthful (Definition 2.14). Moreover, $\mathrm{ATB}$ inherits the error decomposition:
@@ 4.4 Sample Complexity
By Proposition D.2, it suffices to prove the following Rademacher complexity bounds for the function families $F=\{f_{q}\}_{q\in[0,1]}$ and $G=\{g_{q}\}_{q\in[0,1]}$ where $f_{q}(v,y)=(v-y)\mathbb{I}\left[v<q\right]$ and $g_{q}(v,y)=(v-y)\mathbb{I}\left[v\geq q\right]$: for every $(v_{1},y_{1}),\ldots,(v_{T},y_{T})\in[0,1]\times\{0,1\}$,
@@
Now consider the family $H=\{h_{q}\}_{q\in[0,1]}$ where $h_{q}(v,y)=\mathbb{I}\left[v<q\right]$. Clearly, $H$ has VC dimension at most $1$. By Proposition D.5, we have
@@
Observe that $f_{q}(v_{i},y_{i})=\eta_{i}(h_{q}(v_{i},y_{i}))$ for every $i=1,\ldots,T$ and $q\in[0,1]$, where $\eta_{i}$ is the uni-variate function $\eta_{i}(u)=(v_{i}-y_{i})u$ for $u\in\mathbb{R}$. Since $|v_{i}-y_{i}|\leq 1$, the function $\eta_{i}$ is $1$-Lipschitz. Therefore, by Proposition D.3, inequality (40) follows from (42). Inequality (41) can be proved similarly. ∎
@@ 4.5 Computational Efficiency
As we show in the following theorem, $\mathrm{ATB}$ can be computed and approximated efficiently.
Theorem 4.12.
Given $\bm{r},\bm{p}\in[0,1]^{T}$, we can compute $\mathrm{ATB}(\bm{r},\bm{p})$ in time $O(T\log T)$. We can also approximate $\mathrm{ATB}(\bm{r},\bm{p})$ up to arbitrary additive error $\varepsilon>0$ in time $O(T+1/\varepsilon)$.
The algorithm we use to prove Theorem 4.12 is extremely easy to describe and implement. Define
@@
  • $O(T\log T)$ time: sort predictions in increasing order such that $r_{1}\leq r_{2}\leq\dots\leq r_{T}$. Define $r_{0}=0$ and $r_{T+1}=1$.
  • $O(T)$ time: for $q=r_{1},\ldots,r_{T+1}$, calculate $\Delta_{1}(q)$ by scanning predictions in increasing order. Similarly, calculate $\Delta_{2}(q)$ by scanning predictions in decreasing order.
  • $O(T)$ time: Calculate the expectation over threshold $q$: for $t=1,\dots,T+1$, sum up $\Delta_{1}(r_{t})^{2}+\Delta_{2}(r_{t})^{2}$ with weight $|r_{t}-r_{t-1}|$.
If we allow additive error $\varepsilon\in(0,1)$, by Theorem 4.9, we can round the predictions $r_{1},\ldots,r_{T}$ to multiples of $\varepsilon/6$ and then compute $\mathrm{ATB}$ exactly. This makes all predictions $r_{1},\ldots,r_{T}$ lie in a finite set $\{0,\varepsilon/6,2\varepsilon/6,\ldots\}\cap[0,1]$ of size $O(1/\varepsilon)$, so the sorting step can be implemented in time $O(T+1/\varepsilon)$ by bucket sort.
A similar algorithm computes $\ell_{1}\textup{-}\mathrm{ATB}(\bm{r},\bm{p})$ in $O(T\log T)$ time, or approximates $\ell_{1}\textup{-}\mathrm{ATB}(\bm{r},\bm{p})$ up to error $\varepsilon$ in $O(T+1/\varepsilon)$ time. We note that currently known algorithms for computing smCal and $\underline{\textsc{distCal}}$ are much more complicated, with the best known running time being $O(T\log^{2}T)$ and $O(T^{2}\log T)$, respectively, even when $O(1/\sqrt{T})$ additive error is allowed (Hu et al., 2024).
5 Implications of Truthfulness
In this section, we discuss the implications of a calibration measure being truthful. These implications follow from truthfulness alone and do not require soundness or completeness. Aside from the obvious game-theoretic implications by definition, we show that a truthful measure preserves the Blackwell dominance between calibrated predictors. Moreover, Lu et al. (2025), a subsequence work of our paper, show that our construction of a truthful calibration measure mitigates the well-known binning-size selection problem of ECE. Results in this section resonate with the work of Gneiting (2011) which discusses implications of properness (a.k.a., truthfulness) of loss functions.
5.1 Preservation of Blackwell Dominance
One motivation for using a truthful loss function is that a truthful loss correctly preserves the Blackwell dominance among predictors. Consider two calibrated predictors $A$ and $B$. AA Blackwell dominates $B$ if $A$ achieves a lower loss for any proper loss function. We show that, similarly, a truthful calibration measure preserves the Blackwell dominance between calibrated predictors (Theorem 5.2). We also show that recalibrating a miscalibrated predictor must (weakly) reduce its expected error as long as we use a truthful measure (Theorem 5.3). We state both theorems in the binary outcome setting ($y\in\{0,1\}$) for consistency with the rest of the paper, but they generalize naturally to the multi-class setting ($y\in\{1,\ldots,k\}$) with essentially the same proofs.
We say a loss $l:[0,1]\times\{0,1\}\to\mathbb{R}$ is a proper loss if it is interim truthful for eliciting prediction on a size-one sample, i.e., for every true probability $p\in[0,1]$ and every non-truthful report $r\in[0,1]$,
${\operatorname{\mathbb{E}}}_{y\sim\mathsf{Ber}(p)}\left[l(p,y)\right]\leq{\operatorname{\mathbb{E}}}_{y\sim\mathsf{Ber}(p)}\left[l(r,y)\right].$
Throughout this section, we will refer to a proper loss as a truthful error measure evaluating the quality of one prediction, which is different from a calibration measure (or an error measure) evaluating multiple samples.
Blackwell (1951) proposes the Blackwell dominance as the ordering among calibrated (Bayesian) predictors. In statistical decision theory, proper loss functions equivalently evaluate the decision loss when a decision-maker best responds to a prediction, assuming the prediction is correct (see more details in Section B.1). If predictor $A$ Blackwell dominates predictor $B$, AA is uniformly better than $B$ for any decision maker with any arbitrary loss function. This partial order serves as a minimal ranking of trustworthiness among calibrated predictors.
Definition 5.1 (Blackwell Dominance).
Assume predictors $A$ and $B$ are both calibrated. Let $J^{A}$ be the joint distribution of $(r^{A},y)\in[0,1]\times\{0,1\}$ for the prediction of predictor $A$ and the state. Similarly, $J^{B}$ for predictor $B$. AA Blackwell dominates $B$ if for any proper loss function $l$,
${\operatorname{\mathbb{E}}}_{J^{A}}\left[l(r^{A},y)\right]\leq{\operatorname{\mathbb{E}}}_{J^{B}}\left[l(r^{B},y)\right].$
By definition, evaluating by a proper loss preserves the Blackwell dominance among calibrated predictors, which has been one of the main justifications for the use of proper losses (Gneiting, 2011).
We show that this Blackwell preserving property is closed under the composition of samples. Given a truthful error measure on multiple i.i.d. samples from the same predictor, every truthful error measure preserves the same Blackwell dominance.
@@ 5.1 Preservation of Blackwell Dominance
Let Cal be a calibration measure that is ex-ante truthful (Definition 2.12). Let $D$ be an arbitrary joint distribution of $(x,y)\in X\times\{0,1\}$, and let $r^{A},r^{B}:X\to[0,1]$ be two calibrated predictors w.r.t. $D$. Assume that $r^{A}$ Blackwell dominates $r^{B}$. Let $S=\{(x_{t},y_{t})\}_{t\in[T]}$ be a sample of $T$ i.i.d. points drawn from $D$, and let $D_{S}$ denote the (empirical) uniform distribution over $S$. Then
$\operatorname{\mathbb{E}}_{S}[\textsc{Cal}_{D_{S}}(r^{A})]\leq\operatorname{\mathbb{E}}_{S}[\textsc{Cal}_{D_{S}}(r^{B})].$
The proof is deferred to Section B.2.
Theorem 5.2 shows that, for any two calibrated predictors with Blackwell dominance, a truthful error measure preserves the dominance in expectation. Theorem 5.3 furthur characterizes the ordering between calibrated predictors and miscalibrated predictors. Specifically, calibrating a predictor always reduces the expected error measure for a truthful measure.
Theorem 5.3 (Recalibration reduces expected truthful error).
Let Cal be an error measure that is ex-ante truthful (Definition 2.12). Let $J$ be a joint distribution of $(v,y)\in[0,1]\times\{0,1\}$ and define random variable $\widehat{v}:=\operatorname{\mathbb{E}}_{J}[y|v]$ as a function of $v$ (by calibrating $v$). For $T$ i.i.d. examples $(\widehat{v}_{1},v_{1},y_{1}),\ldots,\allowbreak(\widehat{v}_{T},v_{T},y_{T})$, defining $\widehat{\bm{v}}:=(\widehat{v}_{1},\ldots,\widehat{v}_{T}),\bm{v}:=(v_{1},\ldots,v_{T})$ and $\bm{y}=(y_{1},\ldots,y_{T})$, we have
$\operatorname{\mathbb{E}}[\textsc{Cal}(\widehat{\bm{v}},\bm{y})]\leq\operatorname{\mathbb{E}}[\textsc{Cal}(\bm{v},\bm{y})].$
The proof is deferred to Section B.3.
In summary, a truthful error measure preserves the ranking among predictors in the following way.
  • For calibrated predictors, the expected truthful error measure preserves Blackwell ordering.
  • For a miscalibrated predictor, calibrating reduces the truthful error measure in expectation.
Suppose $A$, $B$, and $C$ are predictors such that $B$ is obtained by calibrating $C$, and $A$ Blackwell dominates $B$. With a truthful error measure, the expected error is ordered increasingly by $A$, $B$, and then $C$. As shown in Example 1.1, if evaluated by a non-truthful error measure, both $B$ and $C$ can have a lower (better) error than the dominating predictor $A$, on contrary.
5.2 Robustness to Hyperparameter Selection
One widely observed problem with binned ECE is that the comparison among predictors may flip when the error is estimated with different binning sizes (see, e.g., Minderer et al. 2021). Lu et al. (2025) replicate this non-robustness, which we show in Figure 3(a). This hyperparameter of binning size induces contradicting conclusions about the relationship between the loss (quality) and the calibration error of a predictor.
Subsequent work of ours (Lu et al., 2025) empirically shows that ranking by a truthful calibration measure is robust to hyperparameter selection. A truthful calibration measure preserves the Blackwell ordering among calibrated predictors, which is invariant of the hyperparameter selection. Lu et al. (2025) adopt our truthful Quantile $\ell_{2}\textup{-}\textsc{BinECE}$. They generalize our calibration measure for binary classes to multi-class classification tasks by taking the average over each dimension of prediction.
Lu et al. (2025) explain this robustness as a result of preserving Blackwell dominance, which is invariant of the hyperparameter selection. First, neural network (NN) predictors are approximately calibrated (close to calibration); Second, empirically, there exists a Blackwell dominance between NN predictors across different proper losses. Lastly, for any pairs of (approximately calibrated) predictor with Blackwell dominance, any truthful calibration measure preserves the same ordering.
(a) This plot uses a non-truthful calibration measure, the BinECE. The left estimates ECE with $m=2000$ bins, and more accurate predictors have lower calibration error. The right estimates with $m=20$ bins, and more accurate predictors have higher calibration error.
(b) This plot uses a truthful calibration measure, the Quantile $\ell_{2}\textup{-}\textsc{BinECE}$. The left plot estimates the calibration error with $m=2000$ bins, and the right plot estimates with $m=20$ bins. The ranking between predictors is more consistent for different numbers of bins.
Figure 3: The comparison of calibration measures with different number of bins. Each dot in the plot is a predictor. The $x$-axis plots the log loss (prediction quality), while the $y$-axis plots a calibration error. Figure 3(a) replicates the result in Minderer et al. (2021). The plots are adopted from Lu et al. (2025) with permission.
6 Linear-Time Calibration Tester
@@
In this section, we show that our $\mathrm{ATB}$ and $\ell_{1}\textup{-}\mathrm{ATB}$ are both optimally valid (Definition 2.17) for smCal and $\underline{\textsc{distCal}}$. It is fairly straightforward to show that $\ell_{1}\textup{-}\mathrm{ATB}$ is $O(1/\sqrt{T})$-valid using its constant approximation to smCal (Theorem 4.4) and its sample complexity bound (Theorem 4.10). In Theorem 6.1 below, we show that $\mathrm{ATB}$ is $O(1/\sqrt{T})$-valid as well, and that this is optimal up to constant.
These results imply faster algorithms for solving the calibration testing problem studied by Hu et al. (2024), which requires distinguishing, with large constant success probability, whether a distribution $J$ is perfectly calibrated or has $\underline{\textsc{distCal}}(J)>\varepsilon$ given i.i.d. data points drawn from $J$. This can be solved by computing $\mathrm{ATB}$ or $\ell_{1}\textup{-}\mathrm{ATB}$ on $T=O(1/\varepsilon^{2})$ data points and compare the result with the threshold $\beta_{T}$ in the definition of validity (Definition 2.17). By Theorem 4.12, the running time we need is $O(T\log T)$, which already improves the $O(T\log^{2}T)$ time bound of Hu et al. (2024). Moreover, our Lemmas 6.2 and 6.3 show that it suffices to approximate $\mathrm{ATB}$ up to additive error $1/(2T)$, which can be achieved in time $O(T)$ by Theorem 4.12, giving the first linear-time algorithm for calibration testing.
@@
The proof is completed by choosing $C$ large enough so that $(1/2)(C/3-C^{\prime})^{2}>2$. ∎